OpenStudy (anonymous):

Solve the equation √x^2 + 5 = 3-x

OpenStudy (anonymous):

square both sides to get \[x^2+5=(3-x)^2=x^2-6x+9\] then solve \[x^2+5=x^2-6x+9\] \[5=-6x+9\] \[-6x=-4\] \[x=\frac{2}{3}\]and check the answer